# Autonomy-Aware Clustering: When Local Decisions Supersede Global Prescriptions

### v1

@@ Abstract
Clustering arises in a wide range of problem formulations, yet most existing approaches assume that the entities under clustering are passive and strictly conform to their assigned groups. In reality, entities often exhibit local autonomy, overriding prescribed associations in ways not fully captured by feature representations. Such autonomy can substantially reshape clustering outcomes—altering cluster compositions, geometry, and cardinality—with significant downstream effects on inference and decision-making. We introduce autonomy-aware clustering, a reinforcement (RL) learning framework that learns and accounts for the influence of local autonomy without requiring prior knowledge of its form. Our approach integrates RL with a deterministic annealing (DA) procedure, where, to determine underlying clusters, DA naturally promotes exploration in early stages of annealing and transitions to exploitation later. We also show that the annealing procedure exhibits phase transitions that enable design of efficient annealing schedules. To further enhance adaptability, we propose the Adaptive Distance Estimation Network (ADEN), a transformer-based attention model that learns dependencies between entities and cluster representatives within the RL loop, accommodates variable-sized inputs and outputs, and enables knowledge transfer across diverse problem instances. Empirical results show that our framework closely aligns with underlying data dynamics: even without explicit autonomy models, it achieves solutions close to the ground truth (gap $\sim$3–4%), whereas ignoring autonomy leads to substantially larger gaps ($\sim$35–40%). The code and data are publicly available at https://github.com/salar96/AutonomyAwareClustering.
@@ 2 Related Work
Probabilistic model-based clustering has been widely studied in classical machine learning and applied to diverse domains including image segmentation and topic modeling (Deng & Han, 2018). Fundamental approaches include mixture models such as Gaussian Mixture Models (GMM) and Bernoulli Mixture Models (BMM) (McLachlan & Basford, 1988; McLachlan & Peel, 2000; Figueiredo & Jain, 2002; Zhang et al., 2021), as well as algorithms like Expectation-Maximization (EM) (McLachlan & Krishnan, 2008), probabilistic topic models (Hofmann, 2013) and offline/online Deterministic Annealing (Rose, 1998; Mavridis & Baras, 2022). In such frameworks, data points are not deterministically assigned to clusters; instead, they have soft assignments $\pi_{j|i}\in[0,1]$, with $\sum_{j}\pi_{j|i}=1$ for each data point $i$. These soft assignments naturally accommodate stochasticity in cluster assignment at the level of policy, but they do not capture the local autonomy of the agents where they actively choose or reject assignments.
A line of theoretical work in (Harris et al., 2019; Brubach et al., 2024; Negahbani & Chakrabarty, 2021) have formulated clustering directly under stochastic assignment policy and developed approximation algorithms with provable guarantees. For example,  (Harris et al., 2019) defines a probability distribution, termed as “$k$-lottery,” over possible sets of $K$ centers rather than deterministically selecting a fixed set. The users however, are still passive recipients of assignments, highlighting a gap between these approaches and applications where agents may autonomously accept or reject assignments. In parallel, researchers have cast clustering as a reinforcement learning (RL) problem. One of the earliest examples is Reinforcement Clustering (RC) (Likas, 1999) where each data point’s (assumed passive) assignment to a cluster is treated as an action and the distortion/error acts as the reward signal. More recent work in  (Li et al., 2022; Gowda et al., 2022; Zhu et al., 2025), provide a deep reinforcement learning variant of the approach presented in (Likas, 1999).
@@ 3 Problem Formulation and Solution Methodology
$\pi_{Y}^{\beta}(j|i)=\mathrm{softmax}_{j}\!\big(-\beta\,d_{\mathrm{avg}}(x_{i},y_{j})\big)=\frac{\exp\{-\beta\,d_{\mathrm{avg}}(x_{i},y_{j})\}}{\sum_{\ell=1}^{K}\exp\{-\beta\,d_{\mathrm{avg}}(x_{i},y_{\ell})\}},$ (1)
where $d_{\mathrm{avg}}(x_{i},y_{j}):=\sum_{k=1}^{K}p(k|j,i)\,d(x_{i},y_{k})$ represents the average cost of assigning entity $i$ to cluster $j$ under autonomy-aware reassignments. The Gibbs distribution (1) assigns higher probability to clusters with smaller average costs. The parameter $\beta$ acts as an annealing factor controlling the sharpness of assignments: when $\beta$ is small, $\pi_{Y}^{\beta}(j|i)$ approaches a uniform distribution over clusters (high entropy), encouraging exploration; as $\beta\to\infty$, assignments become increasingly deterministic, converging toward the hard clustering of Problem (P2).
@@ 4 Reinforcement-based method for Autonomy-aware clustering
To account for the local autonomy when it is not known explicitly, we develop a reinforcement-based method to learn the assignment policy $\pi_{Y}^{\beta}(j|i)$, as well as the representative vectors $\{y_{\ell}\}$. Structurally, our proposed learning algorithm parallels Algorithm 1, with the key difference that in the inner while loop (executed at fixed $\beta$), explicit expressions (or updates) of $\pi_{Y}^{\beta}(j|i)$ and $\{y_{\ell}\}$ are replaced with their learning counterparts. We can distinguish between two learning paradigms for autonomy-aware clustering, each motivated by a different reinforcement learning framework: (C1) when $\mathcal{X}$ contains a tractable number of data points, $d(x_{i},y_{k})$ is available in closed form, and $p(l|j,i)$ is independent of $y_{\ell}$}; and (C2) when the dataset is large ($N\gg 1$), $d(x_{i},y_{k})$ is not available in closed form, or the local autonomy depends on $y_{\ell}$}.
@@ Acknowledgments
    Similarly, for the simulations described in Appendix D, the relevant dataset (data point locations and their autonomy levels) is also included in the same supplementary material. The code to reproduce the plots here are straightforward and based on Algorithm 1. The inputs involved are $\beta_{\min}=10^{-3}$, $\beta_{\max}=10^{3}$, $\tau=0.99$, $\rho(i)=\frac{1}{N}$, $\epsilon_{\text{noise}}\sim(10^{-4})\mathcal{N}(0,1)$.
@@ A.1 Case A: When the assumptions in Theorem 1 hold true
    Non-degenerate clusters - there exists $c>0$ such that $p_{\pi_{\rho}}^{Y(t)}(l)\geq c$ for all $l$. In other words, each cluster is a of non-zero mass. Note that this is trivially true for $\pi_{Y}^{\beta}$ in (1) at $\beta<\infty$.
@@
$y_{\ell}=\frac{\sum_{i=1}^{N}\sum_{j=1}^{K}\rho(i)p(l|j,i)\frac{\exp\{-\beta d_{\mathrm{avg}}(x_{i},y_{j})\}}{\sum_{j^{\prime}=1}^{K}\exp\{-\beta d_{\mathrm{avg}}(x_{i},y_{j^{\prime}})\}}x_{i}}{\sum_{i=1}^{N}\sum_{j=1}^{K}\rho(i)p(l|j,i)\frac{\exp\{-\beta d_{\mathrm{avg}}(x_{i},y_{j})\}}{\sum_{j^{\prime}=1}^{K}\exp\{-\beta d_{\mathrm{avg}}(x_{i},y_{j^{\prime}})\}}}.$ (6)
@@
$y_{\ell}(t+1)$ $=\frac{\sum_{i=1}^{N}\sum_{j=1}^{K}\rho(i)p(l|j,i)\frac{\exp\{-\beta d_{\mathrm{avg}}(x_{i},y_{j}(t))\}}{\sum_{j^{\prime}=1}^{K}\exp\{-\beta d_{\mathrm{avg}}(x_{i},y_{j^{\prime}}(t))\}}x_{i}}{\sum_{i=1}^{N}\sum_{j=1}^{K}\rho(i)p(l|j,i)\frac{\exp\{-\beta d_{\mathrm{avg}}(x_{i},y_{j}(t))\}}{\sum_{j^{\prime}=1}^{K}\exp\{-\beta d_{\mathrm{avg}}(x_{i},y_{j^{\prime}}(t))\}}}~\forall~l,$ (7)
$\Rightarrow y_{\ell}(t+1)p_{\pi_{\rho}}^{Y(t)}(l)$ $=\sum_{i=1}^{N}\sum_{j=1}^{K}\rho(i)p(l|j,i)\pi_{Y(t)}^{\beta}(j|i)x_{i},\text{ where}$ (8)
$\pi_{Y(t)}^{\beta}(j|i)=\tfrac{\exp\{-\beta d_{\mathrm{avg}}(x_{i},y_{j}(t))\}}{\sum_{j^{\prime}=1}^{K}\exp\{-\beta d_{\mathrm{avg}}(x_{i},y_{j^{\prime}}(t))\}}$ and $p_{\pi_{\rho}}^{Y(t)}(l)=\sum_{i=1}^{N}\rho(i)\sum_{j=1}^{K}p(l|j,i)\pi_{Y(t)}^{\beta}(j|i)$. Subtracting $p_{\pi_{\rho}}^{Y(t)}(l)y_{\ell}(t)$ from both sides of the equation (8), we obtain
$\big(y_{\ell}(t+1)-y_{\ell}(t)\big)p_{\pi_{\rho}}^{Y(t)}(l)$ $=-\sum_{i=1}^{N}\sum_{j=1}^{K}\rho(i)p(l|j,i)\pi_{Y(t)}^{\beta}(j|i)\big(y_{\ell}(t)-x_{i}\big)$ (9)
$\Rightarrow\big(y_{\ell}(t+1)-y_{\ell}(t)\big)p_{\pi_{\rho}}^{Y(t)}(l)$ $=-\frac{\partial F(Y(t))}{\partial y_{\ell}},$ (10)
@@
Here, $\hat{P}_{\pi_{\rho}}^{Y(t)}=P_{\pi_{\rho}}^{Y}(t)\otimes I_{d}$, $I_{d}$ is a $d\times d$ Identity matrix, $\otimes$ denotes the Kronecker product, $P_{\pi_{\rho}}^{Y(t)}\in\mathbb{R}^{K\times K}$ is a positive definite diagonal matrix with $\big[P_{\pi_{\rho}}^{Y(t)}\big]_{ll}=p_{\pi_{\rho}}^{Y(t)}(l)\geq c>0$, $\nabla F(Y(t))=\Big[\frac{dF(Y(t))}{dy_{1}}^{\top},\cdots,\frac{dF(Y(t))}{dy_{K}}^{\top}\Big]^{\top}\in\mathbb{R}^{Kd}$. At every time instant $t$, we define
@@
$\Rightarrow L(Y(t))\leq\max_{r\in[0,1]}\lambda_{\max}\Big(\big(\hat{P}_{\pi_{\rho}}^{Y(t)}\big)^{-1}\hat{P}_{\pi_{\rho}}^{Y_{r}(t+1)}\Big)=\max_{r\in[0,1]}\max_{1\leq l\leq K}\frac{p_{\pi_{\rho}}^{Y_{r}(t+1)}(l)}{p_{\pi_{\rho}}^{Y(t)}(l)}.$ (26)
@@ Appendix C Sensitivity of $Y$ to the annealing parameter $\beta$
$y_{\ell}=\frac{\sum_{i=1}^{N}p_{\pi_{\rho}}^{Y}(l,i)x_{i}}{p_{\pi_{\rho}}^{Y}(l)},$
where $p_{\pi_{\rho}}^{Y}(l,i)=\sum_{j=1}^{K}\rho(i)\pi_{Y}^{\beta}(j|i)p(l|j,i)$ and $p_{\pi_{\rho}}^{Y}(l)=\sum_{i=1}^{N}p_{\pi_{\rho}}^{Y}(l,i)$. We have that
$\frac{dy_{\ell}}{d\beta}$ $=\frac{1}{p_{\pi_{\rho}}^{Y}(l)}\sum_{i=1}^{N}\frac{dp_{\pi_{\rho}}^{Y}(l,i)x_{i}}{d\beta}-\frac{1}{p_{\pi_{\rho}}^{Y}(l)^{2}}\sum_{i=1}^{N}p_{\pi_{\rho}}^{Y}(l,i)x_{i}\frac{dp_{\pi_{\rho}}^{Y}(l)}{d\beta},$ (50)
where $\dfrac{dp_{\pi_{\rho}^{Y}}(l,i)}{d\beta}=\sum_{j=1}^{K}\rho(i)p(l|j,i)\dfrac{d\pi_{Y}^{\beta}(j|i)}{d\beta}$, where $\pi_{Y}^{\beta}(j|i)$ is the Gibbs’ distribution in (1). We obtain that
@@
$\frac{dy_{\ell}}{d\beta}=\frac{1}{p_{\pi_{\rho}}^{Y}(l)}\sum_{i=1}^{N}\sum_{j=1}^{K}\rho(i)p(l|j,i)\pi_{Y}^{\beta}(j|i)d_{\mathrm{avg}}(x_{i},y_{j})\big(y_{\ell}-x_{i}\big)$
$~~~+\frac{2\beta}{p_{\pi_{\rho}}^{Y}(l)}\sum_{i=1}^{N}\sum_{j=1}^{K}\rho(i)p(l|j,i)\pi_{Y}^{\beta}(j|i)\sum_{k=1}^{K}p(k|j,i)(y_{k}-x_{i})^{\top}\frac{dy_{k}}{d\beta}\big(y_{\ell}-x_{i}\big)$
$~~~-\frac{1}{p_{\pi_{\rho}}^{Y}(l)}\sum_{i=1}^{N}\sum_{j=1}^{K}\rho(i)p(l|j,i)\pi_{Y}^{\beta}(j|i)\sum_{j^{\prime}=1}^{K}\pi_{Y}^{\beta}(j^{\prime}|i)d_{\mathrm{avg}}(x_{i},y_{j^{\prime}})\big(y_{\ell}-x_{i}\big)$
$-\frac{2\beta}{p_{\pi_{\rho}}^{Y}(l)}\sum_{i=1}^{N}\sum_{j=1}^{K}\rho(i)p(l|j,i)\pi_{Y}^{\beta}(j|i)\sum_{j^{\prime}=1}^{K}\pi_{Y}^{\beta}(j^{\prime}|i)\sum_{k=1}^{K}p(k|j^{\prime},i)(y_{k}-x_{i})^{\top}\frac{dy_{k}}{d\beta}(y_{\ell}-x_{i})$ (52)
Multiplying (C) by $p_{\pi_{\rho}}(l)\frac{dy_{\ell}}{d\beta}^{\top}$ on both the sides and summing up over all $l$, we obtain:
$T_{1}:=\sum_{l=1}^{K}p_{\pi_{\rho}}^{Y}(l)\frac{dy_{\ell}}{d\beta}^{\top}\frac{dy_{\ell}}{d\beta}$ $=\sum_{i=1}^{N}\sum_{j,l=1}^{K}\rho(i)p(l|j,i)\pi_{Y}^{\beta}(j|i)d_{\mathrm{avg}}(x_{i},y_{j})\frac{dy_{\ell}}{d\beta}^{\top}(y_{\ell}-x_{i})$
@@
$\qquad-\sum_{i=1}^{N}\sum_{l=1}^{K}p_{\pi_{\rho}}^{Y}(l,i)\sum_{j^{\prime}=1}^{K}\pi_{Y}^{\beta}(j^{\prime}|i)d_{\mathrm{avg}}(x_{i},y_{j^{\prime}})\frac{dy_{\ell}}{d\beta}^{\top}(y_{\ell}-x_{i})$
@@
$~~~\underbrace{\sum_{i=1}^{N}\sum_{j,l,j^{\prime}=1}^{K}\rho(i)p(l|j,i)\pi_{Y}^{\beta}(j|i)\pi_{Y}^{\beta}(j^{\prime}|i)\Big[d_{\mathrm{avg}}(x_{i},y_{j})-d_{\mathrm{avg}}(x_{i},y_{j^{\prime}})\Big]\frac{dy_{\ell}}{d\beta}^{\top}(y_{\ell}-x_{i})}_{=:T_{4}}.$ (54)
@@
$T_{4}$ $\leq\frac{e^{-1}}{\beta}\sum_{i=1}^{N}\sum_{j,l,j^{\prime}=1}^{K}\rho(i)p(l|j,i)\pi_{Y}^{\beta}(j|i)\pi_{Y}^{\beta}(j^{\prime}|i)\frac{dy_{\ell}}{d\beta}^{\top}(y_{\ell}-x_{i})$ (57)
$=\frac{e^{-1}}{\beta}\sum_{i=1}^{N}\sum_{l=1}^{K}p_{\pi_{\rho}}^{Y}(l,i)\frac{dy_{\ell}}{d\beta}^{\top}(y_{\ell}-x_{i})\leq\frac{e^{-1}}{\beta}\sum_{i=1}^{N}\sum_{l=1}^{K}\Big\|\frac{dy_{\ell}}{d\beta}\Big\|R_{\Omega},$ (58)
@@
where $\delta=\min_{l}\big[\lambda_{\min}\big(\frac{\partial^{2}F}{\partial y_{\ell}^{2}}\big)\big]$, and $\lambda_{\min}(\cdot)$ is the minimum eigenvalue. Note that $\sum_{l=1}^{K}\big\|\frac{dy_{\ell}}{d\beta}\big\|\leq\sqrt{K}\big\|\frac{dY}{d\beta}\big\|$ by Cauchy-Schwarz inequality, where

### v3

@@ Abstract
Clustering arises in a wide range of problem formulations, yet most existing approaches assume that the entities under clustering are passive and strictly conform to their assigned groups. In reality, entities often exhibit local autonomy, overriding prescribed associations in ways not fully captured by feature representations. Such autonomy can substantially reshape clustering outcomes—altering cluster compositions, geometry, and cardinality—with significant downstream effects on inference and decision-making. We introduce autonomy-aware clustering, a reinforcement learning (RL) framework that learns and accounts for the influence of local autonomy without requiring prior knowledge of its form. Our approach integrates RL with a Deterministic Annealing (DA) procedure, where, to determine underlying clusters, DA naturally promotes exploration in early stages of annealing and transitions to exploitation later. We also show that the annealing procedure exhibits phase transitions that enable design of efficient annealing schedules. To further enhance adaptability, we propose the Adaptive Distance Estimation Network (ADEN), a transformer-based attention model that learns dependencies between entities and cluster representatives within the RL loop, accommodates variable-sized inputs and outputs, and enables knowledge transfer across diverse problem instances. Empirical results show that our framework closely aligns with underlying data dynamics: even without explicit autonomy models, it achieves solutions close to the ground truth (gap $\sim$3–4%), whereas ignoring autonomy leads to substantially larger gaps ($\sim$35–40%). The code and data are publicly available at https://github.com/salar96/AutonomyAwareClustering.
@@ 2 Related Work
Probabilistic model-based clustering has been widely studied in classical machine learning and applied to diverse domains including image segmentation and topic modeling (Deng & Han, 2018). Fundamental approaches include mixture models such as Gaussian Mixture Models (GMM) and Bernoulli Mixture Models (BMM) (McLachlan & Basford, 1988; McLachlan & Peel, 2000; Figueiredo & Jain, 2002; Zhang et al., 2021), as well as algorithms like Expectation-Maximization (EM) (McLachlan & Krishnan, 2008), probabilistic topic models (Hofmann, 2013) and offline/online Deterministic Annealing (Rose, 1998; Mavridis & Baras, 2022). In such frameworks, data points are not deterministically assigned to clusters; instead, they have soft assignments $\pi_{j|i}\in[0,1]$, with $\sum_{j}\pi_{j|i}=1$ for each $i$. These soft assignments naturally accommodate stochasticity in cluster assignment at the level of policy, but they do not capture the local autonomy of the entities where they actively choose or reject assignments.
A line of theoretical work in (Harris et al., 2019; Brubach et al., 2024; Negahbani & Chakrabarty, 2021) have formulated clustering directly under stochastic assignment policy and developed approximation algorithms with provable guarantees. For example,  (Harris et al., 2019) defines a probability distribution, termed as “$k$-lottery,” over possible sets of $K$ centers rather than deterministically selecting a fixed set. The entities however, are still passive recipients of assignments, highlighting a gap between these approaches and applications where entities may autonomously accept or reject assignments. In parallel, researchers have cast clustering as a reinforcement learning (RL) problem. One of the earliest examples is Reinforcement Clustering (RC) (Likas, 1999) where each entity’s (assumed passive) assignment to a cluster is treated as an action and the distortion/error acts as the reward signal. More recent work in  (Li et al., 2022; Gowda et al., 2022; Zhu et al., 2025), provide a deep reinforcement learning variant of the approach presented in (Likas, 1999).
@@ 3 Problem Formulation and Solution Methodology
$\pi_{Y}^{\beta}(j|i)=\mathrm{softmax}_{j}\!\big(-\beta\,d_{\mathrm{avg}}(x_{i},y_{j})\big)=\frac{\exp\{-\beta\,d_{\mathrm{avg}}(x_{i},y_{j})\}}{\sum_{\ell=1}^{K}\exp\{-\beta\,d_{\mathrm{avg}}(x_{i},y_{\ell})\}}.$ (1)
The Gibbs distribution (1) assigns higher probability to clusters with smaller average costs. The parameter $\beta$ acts as an annealing factor controlling the sharpness of assignments: when $\beta$ is small, $\pi_{Y}^{\beta}(j|i)$ approaches a uniform distribution over clusters (high entropy), encouraging exploration; as $\beta\to\infty$, assignments become increasingly deterministic, converging toward the hard clustering of Problem (P2).
@@ 4 Reinforcement-based method for Autonomy-aware clustering
To account for the local autonomy when it is not known explicitly, we develop a reinforcement-based method to learn the assignment policy $\pi_{Y}^{\beta}(j|i)$, as well as the representative vectors $\{y_{\ell}\}$. Structurally, our proposed learning algorithm parallels Algorithm 1, with the key difference that in the inner while loop (executed at fixed $\beta$), explicit expressions (or updates) of $\pi_{Y}^{\beta}(j|i)$ and $\{y_{\ell}\}$ are replaced with their learning counterparts. We can distinguish between two learning paradigms for autonomy-aware clustering, each motivated by a different reinforcement learning framework: (C1) when $\mathcal{X}$ contains a tractable number of data points, $d(x_{i},y_{k})$ is available in closed form, and $p(\ell|j,i)$ is independent of $y_{\ell}$}; and (C2) when the dataset is large ($N\gg 1$), $d(x_{i},y_{k})$ is not available in closed form, or the local autonomy depends on $y_{\ell}$}.
@@ Acknowledgments
    Similarly, for the simulations described in Appendix D, the relevant dataset (data point locations and their autonomy levels) is also included in the same supplementary material. The code to reproduce the plots here are straightforward and based on Algorithm 1. The inputs involved are $\beta_{\min}=10^{-3}$, $\beta_{\max}=10^{3}$, $\tau=1/0.99$, $\rho(i)=\frac{1}{N}$, $\epsilon_{\text{noise}}\sim(10^{-4})\mathcal{N}(0,1)$.
@@ A.1 Case A: When the assumptions in Theorem 1 hold true
    Non-degenerate clusters - there exists $c>0$ such that $p_{\pi_{\rho}}^{Y(t)}(\ell)\geq c$ for all $\ell$. In other words, each cluster is a of non-zero mass. Note that this is trivially true for $\pi_{Y}^{\beta}$ in (1) at $\beta<\infty$.
@@
$y_{\ell}=\frac{\sum_{i=1}^{N}\sum_{j=1}^{K}\rho(i)p(\ell|j,i)\frac{\exp\{-\beta d_{\mathrm{avg}}(x_{i},y_{j})\}}{\sum_{j^{\prime}=1}^{K}\exp\{-\beta d_{\mathrm{avg}}(x_{i},y_{j^{\prime}})\}}x_{i}}{\sum_{i=1}^{N}\sum_{j=1}^{K}\rho(i)p(\ell|j,i)\frac{\exp\{-\beta d_{\mathrm{avg}}(x_{i},y_{j})\}}{\sum_{j^{\prime}=1}^{K}\exp\{-\beta d_{\mathrm{avg}}(x_{i},y_{j^{\prime}})\}}}.$ (6)
@@
$y_{\ell}(t+1)$ $=\frac{\sum_{i=1}^{N}\sum_{j=1}^{K}\rho(i)p(\ell|j,i)\frac{\exp\{-\beta d_{\mathrm{avg}}(x_{i},y_{j}(t))\}}{\sum_{j^{\prime}=1}^{K}\exp\{-\beta d_{\mathrm{avg}}(x_{i},y_{j^{\prime}}(t))\}}x_{i}}{\sum_{i=1}^{N}\sum_{j=1}^{K}\rho(i)p(\ell|j,i)\frac{\exp\{-\beta d_{\mathrm{avg}}(x_{i},y_{j}(t))\}}{\sum_{j^{\prime}=1}^{K}\exp\{-\beta d_{\mathrm{avg}}(x_{i},y_{j^{\prime}}(t))\}}}~\forall~\ell,$ (7)
$\Rightarrow y_{\ell}(t+1)p_{\pi_{\rho}}^{Y(t)}(\ell)$ $=\sum_{i=1}^{N}\sum_{j=1}^{K}\rho(i)p(\ell|j,i)\pi_{Y(t)}^{\beta}(j|i)x_{i},\text{ where}$ (8)
$\pi_{Y(t)}^{\beta}(j|i)=\tfrac{\exp\{-\beta d_{\mathrm{avg}}(x_{i},y_{j}(t))\}}{\sum_{j^{\prime}=1}^{K}\exp\{-\beta d_{\mathrm{avg}}(x_{i},y_{j^{\prime}}(t))\}}$ and $p_{\pi_{\rho}}^{Y(t)}(\ell)=\sum_{i=1}^{N}\rho(i)\sum_{j=1}^{K}p(\ell|j,i)\pi_{Y(t)}^{\beta}(j|i)$. Subtracting $p_{\pi_{\rho}}^{Y(t)}(\ell)y_{\ell}(t)$ from both sides of the equation (8), we obtain
$\big(y_{\ell}(t+1)-y_{\ell}(t)\big)p_{\pi_{\rho}}^{Y(t)}(\ell)$ $=-\sum_{i=1}^{N}\sum_{j=1}^{K}\rho(i)p(\ell|j,i)\pi_{Y(t)}^{\beta}(j|i)\big(y_{\ell}(t)-x_{i}\big)$ (9)
$\Rightarrow\big(y_{\ell}(t+1)-y_{\ell}(t)\big)p_{\pi_{\rho}}^{Y(t)}(\ell)$ $=-\frac{\partial F(Y(t))}{\partial y_{\ell}},$ (10)
@@
Here, $\hat{P}_{\pi_{\rho}}^{Y(t)}=P_{\pi_{\rho}}^{Y}(t)\otimes I_{d}$, $I_{d}$ is a $d\times d$ Identity matrix, $\otimes$ denotes the Kronecker product, $P_{\pi_{\rho}}^{Y(t)}\in\mathbb{R}^{K\times K}$ is a positive definite diagonal matrix with $\big[P_{\pi_{\rho}}^{Y(t)}\big]_{\ell\ell}=p_{\pi_{\rho}}^{Y(t)}(\ell)\geq c>0$, $\nabla F(Y(t))=\Big[\frac{dF(Y(t))}{dy_{1}}^{\top},\cdots,\frac{dF(Y(t))}{dy_{K}}^{\top}\Big]^{\top}\in\mathbb{R}^{Kd}$. At every time instant $t$, we define
@@
$\Rightarrow L(Y(t))\leq\max_{r\in[0,1]}\lambda_{\max}\Big(\big(\hat{P}_{\pi_{\rho}}^{Y(t)}\big)^{-1}\hat{P}_{\pi_{\rho}}^{Y_{r}(t+1)}\Big)=\max_{r\in[0,1]}\max_{1\leq\ell\leq K}\frac{p_{\pi_{\rho}}^{Y_{r}(t+1)}(\ell)}{p_{\pi_{\rho}}^{Y(t)}(\ell)}.$ (26)
@@ Appendix C Sensitivity of $Y$ to the annealing parameter $\beta$
$y_{\ell}=\frac{\sum_{i=1}^{N}p_{\pi_{\rho}}^{Y}(\ell,i)x_{i}}{p_{\pi_{\rho}}^{Y}(\ell)},$
where $p_{\pi_{\rho}}^{Y}(\ell,i)=\sum_{j=1}^{K}\rho(i)\pi_{Y}^{\beta}(j|i)p(\ell|j,i)$ and $p_{\pi_{\rho}}^{Y}(\ell)=\sum_{i=1}^{N}p_{\pi_{\rho}}^{Y}(\ell,i)$. We have that
$\frac{dy_{\ell}}{d\beta}$ $=\frac{1}{p_{\pi_{\rho}}^{Y}(\ell)}\sum_{i=1}^{N}\frac{dp_{\pi_{\rho}}^{Y}(\ell,i)x_{i}}{d\beta}-\frac{1}{p_{\pi_{\rho}}^{Y}(\ell)^{2}}\sum_{i=1}^{N}p_{\pi_{\rho}}^{Y}(\ell,i)x_{i}\frac{dp_{\pi_{\rho}}^{Y}(\ell)}{d\beta},$ (50)
where $\dfrac{dp_{\pi_{\rho}^{Y}}(\ell,i)}{d\beta}=\sum_{j=1}^{K}\rho(i)p(\ell|j,i)\dfrac{d\pi_{Y}^{\beta}(j|i)}{d\beta}$, where $\pi_{Y}^{\beta}(j|i)$ is the Gibbs’ distribution in (1). We obtain that
@@
$\frac{dy_{\ell}}{d\beta}=\frac{1}{p_{\pi_{\rho}}^{Y}(\ell)}\sum_{i=1}^{N}\sum_{j=1}^{K}\rho(i)p(\ell|j,i)\pi_{Y}^{\beta}(j|i)d_{\mathrm{avg}}(x_{i},y_{j})\big(y_{\ell}-x_{i}\big)$
$~~~+\frac{2\beta}{p_{\pi_{\rho}}^{Y}(l)}\sum_{i=1}^{N}\sum_{j=1}^{K}\rho(i)p(\ell|j,i)\pi_{Y}^{\beta}(j|i)\sum_{k=1}^{K}p(k|j,i)(y_{k}-x_{i})^{\top}\frac{dy_{k}}{d\beta}\big(y_{\ell}-x_{i}\big)$
$~~~-\frac{1}{p_{\pi_{\rho}}^{Y}(\ell)}\sum_{i=1}^{N}\sum_{j=1}^{K}\rho(i)p(\ell|j,i)\pi_{Y}^{\beta}(j|i)\sum_{j^{\prime}=1}^{K}\pi_{Y}^{\beta}(j^{\prime}|i)d_{\mathrm{avg}}(x_{i},y_{j^{\prime}})\big(y_{\ell}-x_{i}\big)$
$-\frac{2\beta}{p_{\pi_{\rho}}^{Y}(\ell)}\sum_{i=1}^{N}\sum_{j=1}^{K}\rho(i)p(\ell|j,i)\pi_{Y}^{\beta}(j|i)\sum_{j^{\prime}=1}^{K}\pi_{Y}^{\beta}(j^{\prime}|i)\sum_{k=1}^{K}p(k|j^{\prime},i)(y_{k}-x_{i})^{\top}\frac{dy_{k}}{d\beta}(y_{\ell}-x_{i})$ (52)
Multiplying (C) by $p_{\pi_{\rho}}(\ell)\frac{dy_{\ell}}{d\beta}^{\top}$ on both the sides and summing up over all $\ell$, we obtain:
$T_{1}:=\sum_{\ell=1}^{K}p_{\pi_{\rho}}^{Y}(\ell)\frac{dy_{\ell}}{d\beta}^{\top}\frac{dy_{\ell}}{d\beta}$ $=\sum_{i=1}^{N}\sum_{j,\ell=1}^{K}\rho(i)p(\ell|j,i)\pi_{Y}^{\beta}(j|i)d_{\mathrm{avg}}(x_{i},y_{j})\frac{dy_{\ell}}{d\beta}^{\top}(y_{\ell}-x_{i})$
@@
$\qquad-\sum_{i=1}^{N}\sum_{\ell=1}^{K}p_{\pi_{\rho}}^{Y}(\ell,i)\sum_{j^{\prime}=1}^{K}\pi_{Y}^{\beta}(j^{\prime}|i)d_{\mathrm{avg}}(x_{i},y_{j^{\prime}})\frac{dy_{\ell}}{d\beta}^{\top}(y_{\ell}-x_{i})$
@@
$~~~\underbrace{\sum_{i=1}^{N}\sum_{j,\ell,j^{\prime}=1}^{K}\rho(i)p(\ell|j,i)\pi_{Y}^{\beta}(j|i)\pi_{Y}^{\beta}(j^{\prime}|i)\Big[d_{\mathrm{avg}}(x_{i},y_{j})-d_{\mathrm{avg}}(x_{i},y_{j^{\prime}})\Big]\frac{dy_{\ell}}{d\beta}^{\top}(y_{\ell}-x_{i})}_{=:T_{4}}.$ (54)
@@
$T_{4}$ $\leq\frac{e^{-1}}{\beta}\sum_{i=1}^{N}\sum_{j,\ell,j^{\prime}=1}^{K}\rho(i)p(\ell|j,i)\pi_{Y}^{\beta}(j|i)\pi_{Y}^{\beta}(j^{\prime}|i)\frac{dy_{\ell}}{d\beta}^{\top}(y_{\ell}-x_{i})$ (57)
$=\frac{e^{-1}}{\beta}\sum_{i=1}^{N}\sum_{\ell=1}^{K}p_{\pi_{\rho}}^{Y}(\ell,i)\frac{dy_{\ell}}{d\beta}^{\top}(y_{\ell}-x_{i})\leq\frac{e^{-1}}{\beta}\sum_{i=1}^{N}\sum_{\ell=1}^{K}\Big\|\frac{dy_{\ell}}{d\beta}\Big\|R_{\Omega},$ (58)
@@
where $\delta=\min_{\ell}\big[\lambda_{\min}\big(\frac{\partial^{2}F}{\partial y_{\ell}^{2}}\big)\big]$, and $\lambda_{\min}(\cdot)$ is the minimum eigenvalue. Note that $\sum_{\ell=1}^{K}\big\|\frac{dy_{\ell}}{d\beta}\big\|\leq\sqrt{K}\big\|\frac{dY}{d\beta}\big\|$ by Cauchy-Schwarz inequality, where